OpenStudy (anonymous):

What is the focus of 64y-4x^2?

OpenStudy (tkhunny):

There is no focus. You have presented an algebraic expression. You will need an actual equation in order to identify a conic section and subsequently identify it's focus.

OpenStudy (anonymous):

Oops my mistake *** The actual equation is 64y = 4x^2

OpenStudy (tkhunny):

That's better. You should have a "Standard Form" for a parabola. What is it? Usually, (y-k) = 4p(x-k)^2 -- or something like that.

OpenStudy (anonymous):

(y-k)2=4p(x-h)

OpenStudy (tkhunny):

Whoops! Of course, I meant the Vertex Form. Do you have that?

OpenStudy (anonymous):

y = a(x – h) 2 + k ?

OpenStudy (tkhunny):

That's it. What does "a" mean in this form?

OpenStudy (anonymous):

Just solved it, but thanks!

OpenStudy (tkhunny):

You're not going to share?

OpenStudy (anonymous):

The answer was actually (0,4) :)